Write an equation of the line containing the specified point and parallel to the indicated line.
(1,4), 3x + y = 5

Answer by lwsshak3(11628)

Write an equation of the line containing the specified point and parallel to the indicated line.
(1,4), 3x + y = 5
**
Standard form of equation for a straight line: y=mx+b, m= slope, b= y-intercept.
For given equation:
3x+y=5
y=-3x+5
slope=-3
For line parallel to given line,
y=-3x+b
slope is the same=-3
solving for b using coordinates of given point (1,4)
4=-3*1+b
b=7
Equation:
y=-3x+7